# A Framework for Fractional Matrix Programming Problems with Applications in FBL MU-MIMO

### v1

@@ VI-B Computational Complexity Comparison
We approximate the number of multiplications to obtain a solution of (60) to compare the computational complexities of our framework and that of the GDA. These algorithms iteratively update the parameters by solving a convex OP. According to [27, Chapter 11], the number of Newton iterations required to solve a convex OP increases with the number of constraints. The GDA updates $\{{\bf W}\}$ by solving (62), having $12K+$ 1 constraints. To solve each Newton iteration, $K$ surrogate functions have to be calculated for the rates, which approximately requires $KN_{BS}^{2}(2N_{BS}+N_{u})$ multiplications. Hence, the approximate number of multiplications to solve (62) is $\mathcal{O}\left[N_{BS}^{2}K\sqrt{2K+1}(2N_{BS}+N_{u})\right]$. Similarly, the number of multiplications to update $\{{\bf W}\}$ in each iteration of our framework can be approximated as $\mathcal{O}\left[N_{BS}^{2}K\sqrt{3K+1}(2N_{BS}+N_{u})\right]$, which is on the same order as the number of multiplications required for the GDA. As the GDA has two loops, the total number of products for solving (60) by the GDA can be limited to $MN$, where $M$ and $N$ are the maximum numbers of iterations for the inner and outer loops, respectively. However, our framework has only one loop, which makes its implementation easier than the GDA and may reduce the total number of iterations.
To sum up, the computational complexity of updating $\{{\bf W}\}$ at each step is of the same order for both algorithms. More specifically, the approximate number of multiplications increases with $K^{3/2}$, $N_{BS}^{3}$, and $N_{u}$ for these algorithms. However, as the GDA is a twin-loop algorithm, the total number of updates for the GDA might be higher than that of our framework.
@@
In Fig. 6, we show the convergence of our algorithm and Dinkelbach-based algorithms, utilizing Monte Carlo simulations. The number of iterations for the Dinkelbach-based algorithms equals the number of updating $\{{\bf W}\}$ to make a fair comparison. Our FMP solver converges faster than Dinkelbach-based algorithms, leading to less computational complexities. The average number of iterations until convergence to a max-min EE in Fig. 6a is $24$ for our framework and $33$ for the GDA. Moreover, our solver requires $76$ iterations on average to solve the maximization of GEE, while the Dinkelbach algorithm needs an average number of $169$ iterations. Note that the convergence of our framework is smoother than Dinkelbach-based algorithms since our solver is a single-loop algorithm. Indeed, the inner loop in Dinkelbach-based algorithms improves the OF much less than the outer loop. Hence, the bigger jumps of Dinkelbach-based algorithms in Fig. 6 correspond to updating the outer loop, making the curves non-smooth.
VII Summary, Conclusion, and Future Works
@@
VII-A Future Research Lines
Our general FMP solver is applicable to diverse MU-MIMO systems such as RIS-aided systems, multiple-access channels, multi-user interference channels, multi-cell BCs, and cognitive radio systems. In this work, we provided a practical example for single-cell BCs operating with and without RIS. In future studies, one can apply the framework to other multi-user systems. Moreover, we studied only a nearly passive diagonal RIS. However, the framework can be used in other RIS architectures, including globally passive [28, 29, 30], non-diagonal RIS [31, 32, 33], active RIS [34], multi-sector [35], simultaneously transmitting and reflecting (STAR) RIS [36, 37, 38] and stacked intelligent surfaces [39]. Additionally, the framework is compatible with other emerging technologies such as fluid antennas.
In addition to the aforementioned technologies, the framework is suitable for complex multiple-access (MA) schemes, utilizing successive interference cancellation (SIC). These MA schemes include, for instance, RSMA [40] and non-orthogonal MA (NOMA) [41]. To this end, the framework in [42] for RSMA and in [43] for NOMA-based schemes can be insightful. Moreover, the framework proposed can also be employed when developing robust designs following the worst-case robustness principle. Additionally, one can apply the FMP solver to scenarios with statistical CSI. This framework also can be used for minimizing overall latency when developing routing schemes for ad hoc systems. Finally, the FMP solver can be harnessed in scenarios with hardware impairment, using improper signaling extending the framework in [44].
@@ Appendix B Proof of Theorem 1
where the notations are defined as in Theorem 1. The constraint in (67b) is not convex even when $f_{mi}(\{{\bf X}\})$ is convex in $\{{\bf X}\}$ since $u_{mi}^{2}$ is convex, instead of being concave. Thus, we leverage the convex-concave procedure (CCP) to convexify (67b) as in (3d), which proves the theorem because the CCP falls into the MM framework [46].
@@ Appendix C Proof of Lemma 1
The OF of (3) is a convex function. Additionally, the constraints in (3b) and (3e) are convex. Hence, when $f_{mi}(\{{\bf X}\})$ and $g_{mi}(\{{\bf X}\})$ are, respectively, non-convex and non-concave, (3) is non-convex due to (3d) and (3c). We can “convexify” these constraints by leveraging the surrogate functions $\tilde{f}_{mi}(\{{\bf X}\})$ and $\tilde{g}_{mi}(\{{\bf X}\})$, obeying the conditions stated in Lemma 1. In this case, we can attain an SP of (1) since all the surrogate functions of (4) satisfy the three conditions in Appendix A.
@@ Appendix G Simulation Parameters
In this appendix, we present the simulation parameters that are chosen based on [47, 18]. More specifically, we assume that the direct channels between the BS and users have a non-line of sight (NLoS) nature, thus following the Rayleigh distribution. Additionally, the channels between the BS and RIS as well as the channels between the RIS and users have a line of sight (LoS) component, hence following the Rician distribution with the Rician factor of $3$. The path-loss exponent is equal to $2.2$ and $3.75$ for the LoS and NLoS channels, respectively, and the large-scale path loss is calculated according to [47, Eq. (59)]. The channel bandwidth is $1.5$ MHz, and the noise power density is $-174$ dBm/Hz. Additionally, we set $N_{{RIS}}=20$, $\epsilon=10^{-5}$ and $n=256$. The other simulation parameters are chosen based on [47, Table II]. Finally, the system topology is depicted in Fig. 7. Indeed, we assume that the users are randomly located in the gray square area of Fig. 7 with a uniform distribution.

### v2

@@ VI-B Computational Complexity Comparison
We approximate the number of multiplications to obtain a solution of (71) to compare the computational complexities of our framework and that of the GDA. These algorithms iteratively update the parameters by solving a convex OP. According to [30, Chapter 11], the number of Newton iterations required to solve a convex OP increases with the number of constraints. The GDA updates $\{{\bf W}\}$ by solving (73), having $12K+$ 1 constraints. To solve each Newton iteration, $K$ surrogate functions have to be calculated for the rates, which approximately requires $KN_{BS}^{2}(2N_{BS}+N_{u})$ multiplications. Hence, the approximate number of multiplications to solve (73) is $\mathcal{O}\left[N_{BS}^{2}K\sqrt{2K+1}(2N_{BS}+N_{u})\right]$. Similarly, the number of multiplications required to update $\{{\bf W}\}$ in each iteration of our framework can be approximated as $\mathcal{O}\left[N_{BS}^{2}K\sqrt{3K+1}(2N_{BS}+N_{u})\right]$, which is on the same order as the number of multiplications required for the GDA. As the GDA involves two loops, the total number of iterations necessitated for solving (71) by the GDA can be limited to $MN$, where $M$ and $N$ are the maximum numbers of iterations for the inner and outer loops, respectively. However, our framework involves only a single loop, which makes its implementation easier than that of the GDA and may reduce the required total number of iterations.
To sum up, the computational complexity of updating $\{{\bf W}\}$ in each step is of the same order for both algorithms. More specifically, the approximate number of multiplications increases with $K^{3/2}$, $N_{BS}^{3}$, and $N_{u}$ for these algorithms. However, as the GDA is a dual-loop algorithm, the total number of updates for the GDA might be higher than for our framework.
@@
In Fig. 6, we show the convergence of our algorithm and Dinkelbach-based algorithms, utilizing Monte Carlo simulations. The number of iterations for the Dinkelbach-based algorithms equals the number of updating $\{{\bf W}\}$ to make a fair comparison. Our FMP solver converges faster than Dinkelbach-based algorithms, leading to less computational complexities. The average number of iterations until convergence to a max-min EE in Fig. 6a is $24$ for our framework and $33$ for the GDA. Moreover, our solver requires $76$ iterations on average to solve the maximization of GEE, while the Dinkelbach algorithm needs an average of $169$ iterations. Note that the convergence of our framework is smoother than that of Dinkelbach-based algorithms, since our solver is a single-loop algorithm. Indeed, the inner loop of Dinkelbach-based algorithms improves the OF much less than the outer loop. Hence, the steps for Dinkelbach-based algorithms seen in Fig. 6 correspond to updating the outer loop, making the curves non-smooth.
VII Extension to Other Practical Scenarios
Our general FMP solver is applicable to diverse MU-MIMO systems such as RIS-aided systems, multiple-access channels, multi-user interference channels (ICs), multi-cell BCs, and cognitive radio systems. In this work, we provide a practical example for single-cell BCs operating with and without RIS, assuming perfect devices and global channel state information (CSI). However, our framework is much more general than the system models and OPs described in Section III, Section IV, and Section V. In this section, we provide insights on how one can apply the general FMP solver to other practical scenarios, specifically in connection with the existing works in the literature. To this end, we do not focus on only MU-MIMO FBL systems.
VII-A Applications in Systems with Hardware Impairments
In this subsection, we elaborate on how our framework applies to multiple-antenna MU systems affected by various sources of hardware imperfections. A common approach for studying hardware impairments (HWI) is to model the aggregated impact of these imperfections as additive Gaussian noise, having a variance proportional to the power of the transmitted/received signal [31, 32, 33, 34, 35]. Here, we consider the scenario examined in one of these references, specifically [35], which derives the rate region of a $K$-user MIMO IC by modeling HWIs as additive Gaussian noise and solving the maximization of the minimum weighted rate as formulated in [35, Eq. (6)]. Instead of the minimum weighted rate OP, our frameworks can be used to solve FMP OPs such as those discussed in Section III and Section IV. In particular, one can employ the concave lower bounds for the user rates developed in [35, Theorem 1] and apply our solution in Section III-B to minimize the sum delay of users in the $K$-user MIMO IC with HWI.
Another common HWI is I/Q imbalance (IQI) [36, 37, 38]. When a device suffers from IQI, its output signal becomes a widely linear transformation of its input signal, rendering the resultant signal improper. In [38], an optimization framework for the $K$-user MIMO IC with IQI was proposed to maximize the weighted sum rate, the weighted minimum rate/EE, and the GEE. By leveraging our FMP solver, the framework in [38] can be extended to tackle the FMP OPs presented in Section III and Section IV. In particular, one can use [38, Lemma 2] to transform the improper signals/noise caused by IQI into real-domain signals. Then, a concave lower bound for the rates can be derived based on the results in [38, Lemma 5]. Once the lower bounds are calculated, Lemma 1 and Lemma 2 can be used to solve the resulting minimization and maximization FMP problems, respectively.
VII-B Applications in Systems with Imperfect CSI
The practical examples considered in the previous sections assume perfect global CSI. However, the proposed framework can also be employed when developing robust designs following the worst-case robustness principle conceived for noisy CSI. Indeed, the results in Section II are general and independent of CSI assumptions. For robust designs following the worst-case robustness principle, each channel is assumed to belong to an uncertainty set with a specified probability [39]. Then system parameters, including beamforming vectors or RIS scattering matrices, are designed to guarantee a specific worst-case performance for all channels within the uncertainty set [39]. To deal with the uncertainty set, one can use the S-procedure or its generalization presented in [40, Lemma 1]. Following this principle, the authors of [41] proposed a robust rate-splitting scheme for maximizing the minimum rate in MISO BCs with FBL. Upon leveraging our FMP frameworks, the scheme in [41] can be extended to solve the FMP OPs of Section III and Section IV in MU-MISO FBL BCs.
VII-C Other Future Research Lines
This subsection outlines additional future research directions for our FMP framework. In the practical examples considered in this paper, we focus on Gaussian signals, although the proposed framework itself is independent of the signaling schemes employed. An interesting research direction is to study OPs obtained for non-Gaussian signals and discrete constellations, which might be even more practical. Moreover, we explain how the framework can be applied to systems with imperfect devices, considering two widely-used HWI models. Another timely research avenue is to investigate the impact of low-resolution analog-to-digital converters (ADCs), especially in applications to ultra-wide-band systems. Finally, we highlight three other promising applications of our FMP framework in the following.
VII-C1 Emerging Technologies for Intelligent Surfaces
In this paper, we study only a nearly passive diagonal RIS. However, the framework can be used in other RIS architectures, including globally passive [42, 43, 44], non-diagonal RIS [45, 46, 47, 48], active RIS [49], multi-sector [50], simultaneously transmitting and reflecting (STAR) RIS [51, 52, 53] and stacked intelligent surfaces [54]. Additionally, the framework is compatible with other emerging technologies such as fluid antennas.
VII-C2 Next Generations of Multiple Access
We consider only treating interference as noise (TIN) in this paper. However, the framework is suitable for complex multiple-access (MA) schemes, utilizing successive interference cancellation (SIC). These MA schemes include, for instance, RSMA [55] and non-orthogonal MA (NOMA) [56]. To this end, the framework in [57] for RSMA and in [29] for NOMA-based schemes can be insightful.
VII-C3 Mobile-edge Computing
Latency minimization is a primary objective in applications related to mobile-edge computing. In these applications, the optimization variables include not only transmission parameters at the transmitters/RIS but also the number of bits offloaded to a computing center via a communication link [22, 23]. Our FMP framework in Section II-A provides a powerful tool for developing latency-aware schemes for MU-MIMO RIS-aided mobile-edge computing systems operating under FBL coding.
VIII Summary and Conclusion
@@ Appendix B Proof of Theorem 1
where the notations are defined as in Theorem 1. The constraint in (78b) is not convex, not even when $f_{mi}(\{{\bf X}\})$ is convex in $\{{\bf X}\}$, since $u_{mi}^{2}$ is convex, instead of being concave. Thus, we employ the concave lower bound in Lemma 11 to convexify (78b) as
$u_{mi}^{2}-f_{mi}(\{{\bf X}\})\geq\\ 2u_{mi}^{(z)}u_{mi}-\left(u_{mi}^{(z)}\right)^{2}-f_{mi}(\{{\bf X}\})\geq 0,\,\forall m,i.$ (79)
This approach is also known as the convex-concave procedure (CCP) and falls into the MM framework [59]. We set $u_{mi}^{(z)}=\sqrt{f_{mi}(\{{\bf X}^{(z)}\})}$ for all $m,i$, since we assume $f_{mi}(\{{\bf X}\})\leq u_{mi}^{2}$, and the bound in (79) is tighter when the equality holds. Upon substituting (78b) with (79) and setting $u_{mi}^{(z)}=\sqrt{f_{mi}(\{{\bf X}^{(z)}\})}$ for all $m,i$, the proof of Theorem 1 is complete.
@@ Appendix C Proof of Lemma 1
Similar to the proof of Theorem 1, we first define variables $u_{mi}\geq 0$ and $t_{mi}>0$, satisfying $f_{mi}(\{{\bf X}\})\leq u_{mi}^{2}$ and $g_{mi}(\{{\bf X}\})\geq t_{mi}$, $\forall m,i$. Hence, (1) can be rewritten as
[…]
$u_{mi}^{2}-f_{mi}(\{{\bf X}\})\!\geq\!0,\,\,\forall m,i,\!\!$ (80b)
$g_{mi}(\{{\bf X}\})\!-\!t_{mi}\geq 0,\,\forall m,i.\!\!$ (80c)
The OF of (80) is a convex function. Additionally, the constraints in (3b) and (3e) are convex. Hence, when $f_{mi}(\{{\bf X}\})$ and $g_{mi}(\{{\bf X}\})$ are, respectively, non-convex and non-concave, the OP in (80) is non-convex due to (80b) and (80c). To “convexify” (80b), one can utilize the concave lower bound in Lemma 11 and the surrogate function $\tilde{f}_{mi}(\{{\bf X}\})$, obeying the conditions stated in Lemma 1, which results in
$u_{mi}^{2}-f_{mi}(\{{\bf X}\})\geq\\ 2u_{mi}^{(z)}u_{mi}-\left(u_{mi}^{(z)}\right)^{2}-\tilde{f}_{mi}(\{{\bf X}\})\geq 0,\,\forall m,i.$ (81)
Note that we set $u_{mi}^{(z)}=\sqrt{f_{mi}(\{{\bf X}^{(z)}\})}$ for all $m,i$, as explained in Appendix B. Additionally, we can convexify (80c) by leveraging the surrogate function $\tilde{g}_{mi}(\{{\bf X}\})$, fulfilling the conditions stated in Lemma 1. In this case, we can attain an SP of (1) since all the surrogate functions of (4) satisfy the three conditions in Appendix A.
@@ Appendix F Useful Inequalities
Lemma 11 ([57]).
For all complex-valued $x_{k}$ and $\bar{x}_{k}$, we have
$\sum_{k}|x_{k}|^{2}\geq\sum_{k}|\bar{x}_{k}|^{2}+2\sum_{k}\mathfrak{R}\left\{\bar{x}_{k}^{*}(x_{k}-\bar{x}_{k})\right\}.$ (93)
@@ Appendix G Simulation Parameters
TABLE VII: List of selected simulation parameters.
Parameter $\tau_{BS-RIS}$ $\tau_{RIS-U}$ $\tau_{BS-U}$ PL0,BS-RIS PL0,RIS-U PL0,BS-U $\iota_{0,BS}$ $\iota_{0,RIS}$ Noise Power D. Ch. Bw. $N_{RIS}$
Value $2.2$ $2.2$ $3.75$ $30$ $30$ $35.9$ $6$ $5.5$ $-174$ dB $1.5$ MHz 20
In this appendix, we present the simulation parameters that are chosen based on [60, 18]. More specifically, we assume that the direct channels between the BS and the users, $\tilde{\mathbf{G}}_{k}$ for all $k$, have a non-line of sight (NLoS) nature, thus following the Rayleigh distribution. It means that each entry of $\tilde{\mathbf{G}}_{k}^{\prime}$ for all $k$ is derived from a zero-mean complex proper Gaussian distribution with unit variance, where we have $\tilde{\mathbf{G}}_{k}=\rho_{k}\tilde{\mathbf{G}}_{k}^{\prime}$, and $\rho_{k}$ is the channel attenuation coefficient due to the large-scale path loss of the direct link between the BS and user $k$. Additionally, the channels between the BS and the RIS, ${\bf G}$, as well as the channels between the RIS and users, ${\bf G}_{k}$ for all $k$, have a line of sight (LoS) component, hence following the Rician distribution. In other words, these channels are given by [26]
${\mathbf{G}}$ $=\sqrt{\frac{\varrho}{1+\varrho}}\tilde{\mathbf{G}}^{\text{LoS}}+\sqrt{\frac{1}{1+\varrho}}\tilde{\mathbf{G}}^{\text{NLoS}},$ (94)
${\mathbf{G}}_{k}$ $=\sqrt{\frac{\varrho}{1+\varrho}}\tilde{\mathbf{G}}^{\text{LoS}}_{k}+\sqrt{\frac{1}{1+\varrho}}\tilde{\mathbf{G}}^{\text{NLoS}}_{k},$ (95)
where $\varrho$ is the Rician factor, which is assumed to be $3$, $\tilde{\mathbf{G}}^{\text{LoS}}$ (or $\tilde{\mathbf{G}}^{\text{LoS}}_{k}$) is the line-of-sight (LoS) component of channel ${\bf G}$ (or ${\bf G}_{k}$), and $\tilde{\mathbf{G}}^{\text{NLoS}}$ (or $\tilde{\mathbf{G}}^{\text{NLoS}}_{k}$) is the non-LoS component of channel ${\bf G}$ (or ${\bf G}_{k}$), which is assumed to follow Rayleigh fading similar to $\tilde{\mathbf{G}}_{k}^{\prime}$. The LoS component $\tilde{\mathbf{G}}^{\text{LoS}}$ is deterministic and given by $\tilde{\mathbf{G}}^{\text{LoS}}=\mathbf{a}_{N_{r}}\left(\psi^{A}\right)\mathbf{a}^{H}_{N_{t}}\left(\psi^{D}\right)$, where $\psi^{A}\sim\text{Unif}[0,2\pi]$ is the angle of arrival, $\psi^{D}\sim\text{Unif}[0,2\pi]$ is the angle of departure, $N_{t}/N_{r}$ is the number of transmit/receive antennas in the corresponding link, and we have
$\mathbf{a}_{N_{r}}\!\left(\psi^{A}\right)$ $\!=\!\!\left[\!1,e^{j\frac{2\pi d\sin(\psi^{A})}{\lambda}},...,e^{j\frac{2(N_{r}-1)\pi d\sin(\psi^{A})}{\lambda}}\right],\!$ (96)
$\!\!\mathbf{a}_{N_{t}}\!\left(\psi^{D}\right)$ $\!=\!\!\left[\!1,e^{j\frac{2\pi d\sin(\psi^{D})}{\lambda}},...,e^{j\frac{2(N_{t}-1)\pi d\sin(\psi^{D})}{\lambda}}\right]\!\!,$ (97)
where $d/\lambda$ is chosen as 1/2 for simplicity [26]. Furthermore, the large-scale path loss in dB is given by [60, Eq. (59)]
$\text{PL}=\text{PL}_{0}+\iota_{0}-10\tau\log_{10}\left(\frac{\varphi}{\varphi_{0}}\right),$ (98)
where $\text{PL}_{0}$ is the path loss at the reference distance $\varphi_{0}=1$m, $\varphi$ is the link distance, $\tau$ is the path-loss exponent, and $\iota_{0}$ is the antenna gain at the transmitter side. Note that the channel attenuation coefficient is $\rho=10^{\text{PL}/10}$.
The system topology is depicted in Fig. 7. We assume that the BS is located at $(0,0,25)$m, where $25$m is the height of the BS. We further assume that the users having a height of $1.5$m are uniformly distributed in the $\times 20$m gray square area of Fig. 7, which is centered at $(130,0,1.5)$m. We also assume that there is only a single RIS with height $15$m, located at $(140,0,15)$m. Furthermore, the simulation parameters are chosen based on [60]. More specifically, the path-loss exponent is equal to $2.2$ and $3.75$ for the LoS and NLoS channels, respectively. Moreover, $\iota_{0}=6$ dB for the BS and $\iota_{0}=5.5$ dB for the RIS. The path loss at the reference distance of $1$m is set to $-35.9$ dB for the direct link between the BS and each user and $-30$ dB for the links related to the RIS. The channel bandwidth is $1.5$ MHz, and the noise power density is $-174$ dBm/Hz. Additionally, we set $N_{{RIS}}=20$, $\epsilon=10^{-5}$ and $n=256$. The simulation parameters are summarized in Table VII.